Two cars (X and Y) leave in the same direction. The car X leaves with a constant speed of 60 km/h and the car Y leaves with a constant speed of 90 km / h. For any distance D (in km) how long does it take (in minutes) for the car Y to reach this distance in relation to the other car.

Answer by TimothyLamb(4379)

s = d/t
t = d/s
---
speed of car y relative to car x:
90 - 60 = 30 kph
---
t = D/30 hr * (60 minutes/hr)
---
answer:
t = 2*D minutes
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---